OpenStudy (anonymous):

Multiple Integrals Anyone? Find the volume of the region bounded on top by the plane z=x+3y+1, on the bottom by the plane z=1/10(x(^2)+y(^2)), and on the sides by the planes x=0, x=3, y=1, y=2.

OpenStudy (perl):

do you know the process

OpenStudy (perl):

first graph the region

OpenStudy (phi):

on the bottom by the plane z=1/10(x(^2)+y(^2)) is a paraboloid shape

OpenStudy (perl):

so thats not a plane, theres a typo in the question

OpenStudy (perl):

Find the volume of the region bounded on top by the plane z=x+3y+1, on the bottom by the *surface* z=1/10(x(^2)+y(^2)), and on the sides by the planes x=0, x=3, y=1, y=2.

OpenStudy (perl):

you can do an idea similiar in one variable calculus, where integrate f(x) - g(x) ,

OpenStudy (anonymous):

I just don't know how to set up the problem. I understand how to integrate with the multiple variables but I don't know what to do with the top and bottom equations.

OpenStudy (perl):

first graph the region over which you will integrate

OpenStudy (perl):

ok took me a while to find your question

OpenStudy (perl):

im trying to get the draw button to work

OpenStudy (anonymous):

So do I subtract the two functions of the top and bottom surfaces? \[\int\limits_{0}^{3}\int\limits_{1}^{2}f(x,y)-g(x,y) dx dy\]

OpenStudy (perl):

you have the limits backwards though

OpenStudy (perl):